OpenStudy (anonymous):

verify the identity and show all work 1- tanxtany= cos(x+y)/cosxcosy

OpenStudy (solomonzelman):

Do I have to work each side independently?

OpenStudy (solomonzelman):

We know that cos(x + y) = cosxcosy - sinxsiny @cicichenxx, right?

OpenStudy (anonymous):

YEAH I THINK SO

OpenStudy (anonymous):

I DONT KNOW HOW TO DO WITH THE LEFT SIDE

OpenStudy (solomonzelman):

cosxcosy - sinxsiny / cosxcosy = cosxcosy/cosxcosy - sinxsiny/cosxcosy = 1 - tanxtany

OpenStudy (anonymous):

should i do something with the left side or do nothing?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

is somebody here?

OpenStudy (solomonzelman):

I was disconnected from the q, sorry.

OpenStudy (anonymous):

i really appreciated your help, but i still confused with the 1-tanxtany

OpenStudy (solomonzelman):

What you do is basically focus on the right side, don't worry about the left side, the right side will be equal to it........ cos(x+y)/cosx cosy (cos x cos y - sin x siny ) / (cos x cos y) cos x cos y/cos x cos y - sin x sin y / cos x cos y 1-tanx tany See?

OpenStudy (anonymous):

yeah i got it thanks and there is another question can you help me with that?

OpenStudy (solomonzelman):

I suck at math, but I can try.

OpenStudy (anonymous):

cot(x+y)=(cotxcoty-1)/(cotx+coty)

OpenStudy (anonymous):

it's ok just try it

OpenStudy (solomonzelman):

Looking at the left side \[\cot(x+y) \] cos(x+y) / sin(x+y) [(cosxcosy - sinxsiny) / (sinxcosy + cosxsiny)] Divide numerator and denominator by sinxsiny, we get (cotxcoty-1)/(coty+cotx),

OpenStudy (anonymous):

yes i got it. and there are two questions i need to do(secx+cscx)/(tanx+cotx)=sinx+cosx cosx/(1-sinx)=secx+tanx

OpenStudy (anonymous):

do you have time?

OpenStudy (solomonzelman):

Yes, but can you please make a new question?

OpenStudy (solomonzelman):

@cicichenxx, tag me if you want to, just like I tagged you!

OpenStudy (anonymous):

how to tag you

OpenStudy (anonymous):

(secx+cscx)/(tanx+cotx)=sinx+cosx

OpenStudy (solomonzelman):

@SolomonZelman, @cicichenxx

OpenStudy (anonymous):

cosx/(1-sinx)=secx+tanx @SolomonZelman

OpenStudy (solomonzelman):

Yup, you just tagged me!

OpenStudy (anonymous):

(secx+cscx)/(tanx+cotx)=sinx+cosx @SolomonZelman

OpenStudy (anonymous):

do you get them?

OpenStudy (anonymous):

i an waiting for your answer

OpenStudy (solomonzelman):

Looking at the left side, sec(x)+csc(x)/tan(x)+cot(x) (1/cos(x) + 1/sin(x)) / (sin(x)/cos(x) + cos(x)/sin(x)) ((sin(x) + cos(x)) / sin(x)cos(x)) / ((sin^2(x) + cos^2(x)) / sin(x)cos(x) sin(x) + cos(x) / 1 sin(x) + cos(x)

OpenStudy (solomonzelman):

Good?

OpenStudy (anonymous):

yeah i am thing about it

OpenStudy (anonymous):

so these questions just do one side is ok?

OpenStudy (solomonzelman):

Yes, you make one side equal to the other. That's the definition of these questions, work one of the sides independently.

OpenStudy (anonymous):

cosx/(1-sinx)=secx+tanx @SolomonZelman

OpenStudy (anonymous):

perhaps this is the lat one i really confusing about those equations

OpenStudy (anonymous):

@SolomonZelman are you here?

OpenStudy (anonymous):

i really confusing about those equations @SolomonZelman

OpenStudy (solomonzelman):

1) cos(x) / (1-sin(x)) Multiply top and bottom by (1 + sin(x)). This gives you [cos(x)(1 + sin(x))] / (1 - sin^2(x)) = [cos(x)(1 + sin(x))] / cos^2(x) = (1 + sin(x)) / cos(x) = sec(x) + tan(x) or Secx + tanx = cosx / (1-sinx) Taking Left side, = Secx + Tanx sec x = 1/cos x and tan x = sin x / cox x Putting the values of secx and tanx you will get = secx + tanx = 1/cosx + sinx / cosx By Taking LCM you will get = (1 + sinx) / cosx Multiply the numerator and denominator by (1-sinx) = (1-sinx)(1+sinx) / cosx(1-sinx) = (1-sin^2 x) / (cos x - sin x cos x) As 1-sin^2 x = cos^2 x therefore; = cos^2 x / [cosx (1-sin x)] Divided numerator and denominator by (cosx) you will get = cosx / (1-sin x)

OpenStudy (anonymous):

i need to think about this, thank for your patient

OpenStudy (solomonzelman):

Anytime!

OpenStudy (solomonzelman):

If you just don;t get it, I can reword it.

OpenStudy (anonymous):

Solve 0.4 cos(x) = sin(2x). SolomonZelman The smallest positive solution is_______ The next smallest positive solution is_____

OpenStudy (anonymous):

Solve 0.4 cos(x) = sin(2x). @SolomonZelman The smallest positive solution is_______ The next smallest positive solution is_____